The length of a rectangle is 3yd longer than width perimeter is 46yd what is the area
Answer by user_dude2008(1862)

2L+2W=46
2(W+3)+2W=46
2W+6+2W=46
4W=40
W=10
L=10+3=13
Area = 10*13 = 130